mathslover (mathslover):

A ball is released from the top of a tower of height h meters. It takes T second to reach the ground. What is the position f the ball in T/3 second

mathslover (mathslover):

I did like this : \[\large{T =\sqrt{\frac{2h}{g}}}\] \[\large{v^2=u^2+2gh'}\] \[\large{g^2(T^2)=2gh'}\] \[\large{g^2*\frac{2h}{g}*\frac{1}{9}=2gh'}\] \[\large{\frac{h}{9}=h'}\]

OpenStudy (anonymous):

we can directly do it like this also rit....h=1/2*g*T/3*T/3===>h/9

mathslover (mathslover):

but the answer is 8h / 9

OpenStudy (ajprincess):

When u measure from the top h'=h/9 . it is h-h/9=8h/9 from the ground.

OpenStudy (ash2326):

@mathslover In your coordinate system you took the top of tower as 0 and the ground as h so when t=T/3 the ball is at h/9 from tower so, it'd be 8h/9 from the ground

mathslover (mathslover):

hmn thnks to both @ajprincess and @ash2326 .. i got the mistake now

OpenStudy (anonymous):

tnx guys^ ^

mathslover (mathslover):

:D this is good , that you also learnt something from this thanks to all

OpenStudy (ajprincess):

U r welcome.:D

mathslover (mathslover):

:)